. Derive Schrodinger wave equation from Cartesian coordinates into Spherical polar coordinates. 1 π‘Ÿ 2 πœ• πœ•π‘Ÿ (π‘Ÿ 2 πœ•πœ“ πœ•π‘Ÿ ) + 1 π‘Ÿ 2π‘ π‘–π‘›πœƒ πœ• πœ•πœƒ (π‘ π‘–π‘›πœƒ πœ•πœ“ πœ•πœƒ) + 1 π‘Ÿ 2𝑠𝑖𝑛2πœƒ πœ• 2πœ“ πœ•πœ™2 + 2π‘š ℏ 2 (𝐸 βˆ’ π‘ˆ)πœ“ = 0


1
Expert's answer
2021-01-14T10:40:36-0500

Let's first write (stationary) Schrodinger equation in cartesian coordinates :

"-\frac{\hbar^2}{2m} (\frac{\partial^2}{\partial x^2}+\frac{\partial^2}{\partial y^2}+\frac{\partial^2}{\partial z^2})\psi +V\psi = E\psi"

"-\frac{\hbar^2}{2m} \Delta\psi + (V-E)\psi=0" , where "\Delta" is Laplace operator

"\Delta\psi + \frac{2m}{\hbar^2} (E-V)\psi=0"

Now it is just enough to use the expression of "\Delta" in spherical coordinates :

"\frac{1}{r^2} \frac{\partial}{\partial r}(r^2\frac{\partial \psi}{\partial r}) + \frac{1}{r^2 \sin\theta} \frac{\partial}{\partial\theta} (\sin\theta \frac{\partial\psi}{\partial\theta}) + \frac{1}{r^2\sin^2\theta} \frac{\partial^2\psi}{\partial\phi^2}+\frac{2m}{\hbar^2}\psi =0"

The expression of "\Delta" in spherical coordinates can be found, for example, in Wikipedia : Laplace operator - Wikipedia . We can also, of course, calculate it directly :

"\frac{\partial}{\partial x} = \frac{\partial r}{\partial x}\frac{\partial}{\partial r} + \frac{\partial \theta}{\partial x}\frac{\partial}{\partial \theta} + \frac{\partial \phi}{\partial x}\frac{\partial}{\partial \phi}" (chain rule)

"\frac{\partial}{\partial x} = \sin\theta \cos\phi \frac{\partial}{\partial r} -\frac{\cos\theta \cos\phi}{r}\frac{\partial}{\partial\theta}-\frac{\sin\phi}{r\sin\theta} \frac{\partial}{\partial\phi}" (using the expressions of spherical coordinates in cartesian coordinates)

The same calculation for "y, z" gives :

"\frac{\partial}{\partial y} = \sin\theta \sin\phi \frac{\partial}{\partial r} -\frac{\cos\theta \sin\phi}{r}\frac{\partial}{\partial\theta} + \frac{\cos\phi}{r\sin\theta} \frac{\partial}{\partial\phi}"

"\frac{\partial}{\partial z} = \cos\theta \frac{\partial}{\partial r} -\frac{\sin\theta}{r}\frac{\partial}{\partial\theta}"

And now we find by direct calculation (as "\frac{\partial^2}{\partial x^2} = (\frac{\partial}{\partial x})^2" and same for other coordinates) :

"\Delta = \frac{1}{r^2} \frac{\partial}{\partial r}(r^2\frac{\partial}{\partial r}) + \frac{1}{r^2 \sin\theta} \frac{\partial}{\partial\theta} (\sin\theta \frac{\partial}{\partial\theta}) + \frac{1}{r^2\sin^2\theta} \frac{\partial^2}{\partial\phi^2}"